# Learning Encodings by Maximizing State Distinguishability: Variational Quantum Error Correction

### v1

@@ 1 Introduction
[…]
Figure 1: We propose a procedure that tailors quantum error correction encodings to arbitrary noise channels. In our setup, a pure state |$\ket{\psi}$ is encoded into a larger ancilla system, and a noise channel then acts on the involved wires. We illustrate this with pairs of states |$\ket{\psi_{0}}=\ket{0}$ and |$\ket{\psi_{1}}=\ket{+}$, which in the noise-free case can be visualized on the top and right side of a unit sphere–the Bloch sphere in the single-qubit setup. Under noise, e.g., a combination of depolarizing and amplitude damping processes, the sphere shrinks and the states are drawn inwards. We quantify this by computing the trace distance between the original states (yellow line) and subtracting the trace distance between their noisy counterparts (orange line). To optimize the encoding for arbitrary states, we minimize the average distinguishability loss over pairs of states drawn from a two-design, thereby encouraging minimal sphere shrinkage (shown by blue arrows) and enabling high-fidelity recovery. For proof of principle, we implement this procedure as a variational quantum algorithm using a parameterized encoding ansatz, calling it variational quantum error correction (VarQEC).
@@
    We incorporate our approach into a variational quantum algorithm, termed variational quantum error correction. This also integrates elements from the QVECTOR approach [46] (see Section 2) for learning recovery operations. We provide an end-to-end implementation to encourage reproducibility and future extensions.
@@
The remainder of this paper is organized as follows: In Section 2, we position our work within the literature on AI-enhanced QEC. Section 3 summarizes the background on noise channels, distance measures for quantum states, and QEC. In Section 4, we formally introduce the distinguishability loss, formulate the VarQEC algorithm, and derive theoretical properties of the learned codes. The empirical results in Section 5 demonstrate the effectiveness of the VarQEC codes compared to standard QEC codes. In Section 6, we deploy code instances on IBM and IQM hardware systems to underline the practical relevance of our procedure. We identify current limitations of the VarQEC procedure and outline future improvements and extensions in Section 7. Finally, in Section 8, we provide an overview and discuss our results in the context of QEC and eFTQC.
@@ 2 Related Work
Classical Machine Learning for QEC. The integration of AI into QEC began around 2018 with the pioneering work of Fösel et al. [31], which employed reinforcement learning (RL). In this approach, a reward function was constructed from several components, including the recoverable quantum information, protection reward, and recovery reward. This methodology was used to train encoding and recovery circuits on systems with up to four physical qubits and one logical qubit, primarily focusing on bit-flip noise. While the recoverable quantum information is loosely related to our proposed distinguishability loss, both being based on the trace distance, the latter is much more suitable as an easy-to-evaluate loss function, as we briefly discuss in Section 4.1.
@@
A modified RL routine for QEC has recently been proposed, formulating reward functions based on the Knill-Laflamme conditions [71, 72]. Aspects of fault tolerance have also been incorporated by learning flag gadgets [104]. RL has further been studied for modifying and combining existing codes [69, 61], autonomous QEC [105, 74], and the discovery of low-weight quantum low-density parity-check (qLDPC) codes [40]. Additionally, RL has been applied to the discovery of logical gates within specific experimental setups [44]. Alternative ML methods have also been explored for various QEC scenarios, including supervised and unsupervised learning techniques [8, 99, 106, 59], as well as evolutionary approaches [91, 102].
@@
Variational Techniques for QEC. Variational methods have gained attention in the context of QEC due to their flexibility and adaptability to quantum hardware. The first proposal to formulate QEC as a variational algorithm was proposed in 2017 by Johnson et al. [46], dubbed QVECTOR. In this approach, the fidelity after recovery is used as a loss function to simultaneously train the free parameters of encoding and recovery circuits. The authors demonstrate the effectiveness of the approach in low-probability amplitude and phase damping noise scenarios, using systems with up to five physical qubits and a single logical qubit. We combine our proposal of learning encoding circuits with parts of the QVECTOR concept to generate recovery operations to demonstrate a full QEC cycle. Details are provided in Section 5.3. A modification to the QVECTOR approach proposes to enhance trainability by employing a Wasserstein version of the fidelity [108]. However, in our evaluations, this improvement did not manifest.
@@
Other works have investigated the use of convolutional quantum neural networks [25] and quantum autoencoders for QEC [60, 80]. In some instances, these methods function more as denoising algorithms than as active QEC strategies. Different techniques aim to generate codes by formulating the task of QEC as finding the ground state of a Hamiltonian [103, 100, 18]. However, these approaches assume full direct access to the error operators and stabilizers. Additionally, preliminary work has been conducted on finding logical operations for given codes via variational techniques [23].
@@ 3 Background
As mentioned before, quantum systems are highly susceptible to errors due to decoherence and interactions with the environment [96]. To build reliable quantum computers, it is crucial to understand how quantum information is affected by noise and how to correct for the introduced errors. In this section, we introduce the underlying concepts on noise channels, distance measures for quantum states, and error correction that go into the VarQEC procedure we introduce in this paper.
@@ 3.1 Quantum Noise Channels
Quantum noise channels are mathematical models that describe how quantum states evolve in the presence of noise and decoherence. A pure $n$-qubit quantum state is typically described using Dirac notation as |$\ket{\psi}\in\mathcal{H}^{\otimes n}$, where $\mathcal{H}$ denotes the Hilbert space, and it holds $\expectationvalue{\psi|\psi}=1$. Using the density matrix formalism, one can express the same state as $\rho=\ket{\psi}\bra{\psi}$ |, which is of rank one for pure states. More general mixed states represent a statistical ensemble of pure states and can be described by a density matrix
$\rho=\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}},$ (1)
@@
where $p$ is the depolarization probability and $X$, $Y$, and $Z$ are the Pauli operators representing bit-flip, bit-flip and phase-flip, and phase-flip errors, respectively [70]. Arbitrary quantum errors in multi-qubit systems can be expanded as linear combinations of tensor products of Pauli operators. Consequently, correcting bit-flip and phase-flip errors on each qubit inherently addresses more general multi-qubit error scenarios [35]. In many practical scenarios, and most of this paper, errors are assumed to act independently on each qubit, leading to a composite noise model
@@
where $\mathcal{N}_{j}$ is a noise channel on qubit $j$. Considerations and experiments on correlated noise are conducted in Section 5.2 and a more detailed discussion can be found in Section E.4. A more rigorous and extensive definition of the quantum noise channels considered in this work can be found in Section A.2.
@@ 3.2 Distance Measures for Quantum States
Quantifying the difference between quantum states is essential for the task of quantum error correction. Measures of state similarity or distinguishability help in evaluating the performance of quantum algorithms and error correction codes [70]. A more extensive introduction and discussion on relations between the metrics introduced in the following can be found in Section A.1.
@@
where $\|A\|_{1}=\mathrm{Tr}\sqrt{A^{t}A}$ denotes the trace norm of the operator $A$ [33]. The trace distance quantifies the distinguishability between quantum states and is bounded between 00 and $1$, with 00 indicating identical states and $1$ representing completely distinguishable (orthogonal) states. These bounds and operational meaning, as well as the fact that it defines a full metric, contribute to the definition of a stable loss function in Section 4. The trace distance is invariant under unitary transformations, i.e.,
@@ 3.3 Quantum Error Correction
The first and crucial step of such an error correction procedure is to find an encoding operation $U_{\mathrm{enc}}$ that maps the $k$-qubit pure state into a larger composite $n$-qubit system, creating redundancy that can protect against errors [14]. The encoding operation produces the encoded state using the original state $\rho$ and ancilla qubits initialized to |$\ket{0}$:
$\rho_{L}=U_{\mathrm{enc}}\left(\rho\otimes\ket{0}\bra{0}^{\otimes n-k}\right)U_{\mathrm{enc}}^{\dagger}$ (9)
After the state passes through some noise channel $\mathcal{N}$, the resulting state is mixed and can be described as
@@
$\hat{\rho}_{L}=\mathrm{Tr}_{r}\left(U_{\text{rec}}\left(\tilde{\rho}_{L}\otimes\ket{0}\bra{0}^{\otimes r}\right)U_{\text{rec}}^{\dagger}\right)$ (11)
@@
with $\delta$ being a small positive error threshold [10]. In the experimental section, we discuss that some instances of the learned code reduce $\delta$ to practically 00, allowing for exact correction. An extensive discussion of our work in the context of exact and approximate error correction can be found in Section B.3.
A QEC code is often characterized by its code parameters [[$n,k,d$]], where $n$ is the number of physical qubits in the encoding, $k$ is the number of encoded logical qubits, and $d$ is the code distance [15]. This distance is the minimum number of qubits that need to affected by an error operation to transform one valid codeword into another one. In other words, it is the minimum number of non-trivial single-qubit Pauli errors that break the code. A distance-$d$ QEC code can detect up to $d-1$ errors and correct up to $\lfloor\frac{d-1}{2}\rfloor$. Therefore, a larger code distance implies a higher capability to correct errors. However, in general, a larger code distance requires a larger number of physical qubits, which is also quantified by the code rate $\frac{k}{n}$.
In practice, one needs to frequently conduct syndrome extraction and recovery
@@ 4 Learning Error Correction Codes
In this section, we introduce our method for learning QEC codes by maximizing state distinguishability. We define a loss function based on the trace distance between quantum states, realize the method as a variational quantum algorithm, discuss how to perform error correction using recovery operations, and analyze the properties of the learned codes.
$\mathcal{N}$|$\ket{0}$$\ket{+}$|$\ket{\tilde{0}}$$\ket{\tilde{+}}$
(a) Effect of a noise channel on a single physical qubit visualized on the Bloch sphere. The shrinking of the sphere and the shift of the center towards the |$\ket{0}$-state can, e.g. be caused by depolarizing combined plus amplitude damping noise [38]. The blue line marks the distance of optimal |$\ket{0}$ and $\ket{+}$ states, the red one marks the distance after the noise channel. For the single-qubit setup, one can equivalently use the trace or the Euclidean distance, yielding $T(\ket{0},\ket{+})=\frac{1}{\sqrt{2}}$.
$\mathcal{N}^{\otimes}$|$\ket{0}_{L}$|$\ket{+}_{L}$|$\ket{\tilde{0}}_{L}$|+~$\ket{\tilde{+}}_{L}$
(b) Effect of a composite noise channel on an encoded multi-qubit system. While these are not Bloch spheres, for the purpose of visualizing the trace distance, the undistorted sphere still has unit length. This follows from the invariance of the trace distance under unitary operations, consequently $T(\ket{0}_{L},\ket{1}_{L})=\frac{1}{\sqrt{2}}$.
$\mathcal{N}^{\otimes}$|$\ket{0}_{L}$|$\ket{+}_{L}$|$\ket{\tilde{0}}_{L}$|+~$\ket{\tilde{+}}_{L}$
@@
Figure 2: Effect of noise on a physical state (a) and an encoded logical state (b). Under the assumption of a sophisticated encoding operation we can satisfy $T(\ket{\tilde{0}},\ket{\tilde{1}})<T(\ket{\tilde{0}}_{L},\ket{\tilde{1}}_{L})$. This allows us to reduce the lost trace distance $\Delta_{T}(\ket{0},\ket{1})$. In practice, we have to minimize this measure over all possible state pairs, or alternatively over all combinations of the spherical two-design in (c).
@@ 4.1 Lost Trace and Distinguishability Loss
where $\rho_{L}$ and $\sigma_{L}$ are prepared using the potentially parameterized encoding operation $U_{\mathrm{enc}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\Theta)}$ as described in Eq. 9. Due to the trace distance’s invariance under both tensoring with pure states and unitary transformations, we can also express the first term as $T(\rho_{L},\sigma_{L})$, but the former is usually more straightforward to compute. This measure is also indicated in Fig. 1 and its components are explicitly marked in Fig. 2.
@@
For practical optimization, evaluating the worst-case loss is challenging because it involves a maximization over an infinite set of state pairs. In principle, it is possible to use techniques from semi-definite programming [88], but it is hard to obtain guarantees for the efficiency of finding the solution. Alternatively, under certain restrictions on ansätze and noise models, one can try to come up with efficient optimization-based evaluation methods, as e.g. proposed for the recoverable quantum information [31]. However, we empirically found that the worst-case formulation also has the drawback that employing it as a loss function leads to severe instabilities in the training procedure, which aligns with findings from classical machine learning [39, 7]. To address this, we define the average-case distinguishability loss as
[…]
@@ 4.2 Variational Quantum Error Correction
Figure 3: Pipeline of learning VarQEC encodings with the distinguishability loss. Pairs of $k$-qubit states from a two-design $\mathcal{S}$ are prepared and encoded into an $n$-qubit Hilbert space with the encoding unitary $U_{\text{enc}}(\Theta)$. The resulting encoded states are subjected to the noise channel $\mathcal{N}$, before the trace distance between the disturbed states is evaluated. The difference to the baseline trace distance is computed and denoted as the lost trace distance $\Delta_{T}(\rho,\sigma;\mathcal{N};\Theta)$. This information is averaged over all state pairs $\rho,\sigma\in\mathcal{S}$, giving the two-design approximation to the average-case distinguishability loss $\mathcal{D}_{\mathcal{S}}(\mathcal{N};\Theta)$. This objective is minimized with a suitable technique, in this paper instantiated as a variational quantum algorithm utilizing gradient-based optimization. The converged encoding constitutes a [[$n,k$]] code – either for exact or approximate QEC, depending on the setup – finetuned for the noise channel $\mathcal{N}$.
In this work, for a proof-of-principle demonstration of the concept, we choose to formulate the task as a variational quantum algorithm [19]. We first select a variational ansatz $U_{\mathrm{enc}}(\Theta)$ for the encoding unitary, with trainable parameters $\Theta$. Therefore, we utilize a newly developed randomized entangling ansatz, detailed in Section D.1. The parameters can be optimized using either gradient-free or gradient-based optimization methods. We employ the quasi-Newton limited-memory BFGS (L-BFGS) algorithm [58] due to its fast convergence for medium-sized models. However, this choice is not mandatory, and other optimization techniques like standard gradient descent can also be used. To enhance performance, in future work, one can also focus on the choice of optimal circuit architectures, using e.g. differentiable architecture search [107] or incrementally growing the ansatz [37].
@@
Due to its resemblance to a variational quantum algorithm, we denote this realization of the training procedure as variational quantum error correction. The resulting circuits are dubbed VarQEC codes, and grouped by their code parameters [[$n,k$]]. While in Section 4.4 we also derive a methodology to quantify the code distance $d$ of VarQEC codes, in general, and by construction, the VarQEC procedure does not explicitly enforce the ability to recover information under specific error terms. Rather, the distinguishability loss is used to automatically focus on specific error terms, to minimize the overall loss of information. The described training procedure is conceptually illustrated in Fig. 3.

### v2

@@ Abstract
keywords:
quantum computing, machine learning, quantum error correction, variational quantum algorithm
@@ 1 Introduction
    We incorporate our approach into a variational quantum algorithm, termed variational quantum error correction. This also integrates elements from the QVECTOR approach [50] for learning recovery operations. We provide an end-to-end implementation to encourage reproducibility and future extensions.
@@
The remainder of this paper is organized as follows: In section˜2, we position our work within the literature on AI-enhanced QEC. Section˜3 summarizes the background on noise channels, distance measures for quantum states, and QEC. In section˜4, we formally introduce the distinguishability loss, formulate the VarQEC algorithm, and derive theoretical properties of the learned codes. The empirical results in section˜5 demonstrate the effectiveness of the VarQEC codes compared to standard QEC codes. In section˜6, we deploy code instances on IBM and IQM hardware systems to underline the practical relevance of our procedure. We identify current limitations of the VarQEC procedure and outline future improvements and extensions in section˜7. Finally, in section˜8, we provide an overview and discuss our results in the context of QEC and eFTQC.
@@ 2 Related Work
Classical Machine Learning for QEC. The integration of AI into QEC began around 2018 with the pioneering work of Fösel et al. [34], which employed reinforcement learning (RL). In this approach, a reward function was constructed from several components, including the recoverable quantum information, protection reward, and recovery reward. This methodology was used to train encoding and recovery circuits on systems with up to four physical qubits and one logical qubit, primarily focusing on bit-flip noise. This notion of recoverable quantum information is related to our proposed distinguishability loss, both being based on the trace distance. However, we argue in section˜4.1 that only the latter constitutes a loss function that is efficient to evaluate for complex noise structures. Furthermore, while the recoverable quantum information was restricted to stabilizer codes, our formulation extends to non-additive codes.
@@
A modified RL routine for QEC has recently been proposed, formulating reward functions based on the Knill-Laflamme conditions [78, 77]. Aspects of fault tolerance have also been incorporated by learning flag gadgets [114]. RL has further been studied for modifying and combining existing codes [75, 66, 98, 112], autonomous QEC [116, 113, 80], and the discovery of gadgets for low-density parity-check (qLDPC) codes [44, 33]. Additionally, RL has been applied to the discovery of logical gates within specific experimental setups [48]. Other QEC scenarios have also been considered through the lens of techniques from supervised and unsupervised learning techniques [8, 107, 115, 64], evolutionary approaches [99, 110], as well as various other ML approaches [41, 91, 26].
@@
Variational Techniques for QEC. Variational methods have gained attention in the context of QEC due to their flexibility and adaptability to quantum hardware. The first proposal to formulate QEC as a variational algorithm was proposed in 2017 by Johnson et al. [50], dubbed QVECTOR. In this approach, the fidelity after recovery is used as a loss function to simultaneously train the free parameters of encoding and recovery circuits. The authors demonstrate the effectiveness of the approach in low-probability amplitude and phase damping noise scenarios, using systems with up to five physical qubits and a single logical qubit. However, we observed that the concept of simultaneously learning encoding and recovery is limited to rather small code instances and simple low-probability noise structures, due to training not converging in more complex scenarios. A modification to the QVECTOR approach proposes to enhance trainability by employing a Wasserstein version of the fidelity [118]. However, in our evaluations, this improvement did not manifest. In contrast to QVECTOR, our proposed technique modularizes the training of the encoding and recovery operations, allowing for more flexibility and the consideration of complex noise structures.
@@
Other works have investigated the use of convolutional quantum neural networks [27] and quantum autoencoders for QEC [65, 86, 61]. In some instances, these methods function more as denoising algorithms than as active QEC strategies. Different techniques aim to generate codes by formulating the task of QEC as finding the ground state of a Hamiltonian [111, 108, 19]. However, these approaches assume full direct access to the error operators and stabilizers. Additionally, preliminary work has been conducted on finding logical operations for given codes via variational techniques [24].
@@ 3 Background
As mentioned before, quantum systems are highly susceptible to errors due to decoherence and interactions with the environment [104]. To build reliable quantum computers, it is crucial to understand how quantum information is affected by noise and how to correct for the introduced errors. In this section, we introduce the underlying concepts on noise channels, distance measures for quantum states, and error correction that go into the VarQEC procedure we introduce in this paper. Supplementary details can be found in Appendix˜A.
@@ 3.1 Quantum Noise Channels
[…]
$|,\displaystyle\rho=\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}$, (1)
@@
where $p$ is the depolarization probability and $X$, $Y$, and $Z$ are the Pauli operators representing bit-flip, bit-flip and phase-flip, and phase-flip errors, respectively [76]. General single-qubit Pauli noise can be described as
$\mathcal{N}(\rho)=(1-p)\rho+p_{X}X\rho X^{\dagger}+p_{Y}Y\rho Y^{\dagger}+p_{Z}Z\rho Z^{\dagger},$ (4)
with overall nosie strength $p=p_{X}+p_{Y}+p_{Z}$. Arbitrary quantum errors in multi-qubit systems can be expanded as linear combinations of tensor products of Pauli operators. Consequently, correcting bit-flip and phase-flip errors on each qubit inherently addresses more general multi-qubit error scenarios [38]. In the context of stabilizer-based QEC, this can be understood as syndrome measurements effectively projecting general errors into Pauli errors [55].
In many practical scenarios, and most of this paper, errors are assumed to act independently on each qubit, leading to a composite noise model
@@
where $\mathcal{N}_{j}$ is a noise channel on qubit $j$. Considerations and experiments on correlated noise are conducted in section˜5.2.
@@ 3.2 Distance Measures for Quantum States
Quantifying the difference between quantum states is essential for the task of quantum error correction. Measures of state similarity or distinguishability help in evaluating the performance of quantum algorithms and error correction codes [76].
@@
where $\|A\|_{1}=\mathrm{Tr}\sqrt{A^{t}A}$ denotes the trace norm of the operator $A$ [36]. The trace distance quantifies the distinguishability between quantum states and is bounded between 0 and $1$, with 0 indicating identical states and $1$ representing completely distinguishable (orthogonal) states. These bounds and operational meaning, as well as the fact that it defines a full metric, contribute to the definition of a stable loss function in section˜4. The trace distance is invariant under unitary transformations, i.e.,
@@ 3.3 Quantum Error Correction
[…]
The After the state passes through some noise channel $\mathcal{N}$, the resulting state is mixed and can be described as
@@
[…]
@@
with $\delta$ being a small positive error threshold [10]. In the experimental section, we discuss that some instances of the learned code reduce $\delta$ to practically 0, allowing for exact correction.
An arbitrary QEC code can be characterized by its code parameters
$((n,K=2^{k},d)),$ (15)
where $n$ is the number of physical qubits to encode $k$ logical qubits (general non-additive codes do not restrict $k=\log_{2}K$ to a whole number, but in this work we do not consider such codes). Furthermore, $d$ is the code distance [15], which gives the minimum number of qubits that need to be affected by an error operation to transform one valid codeword into another one. In other words, it is the minimum number of non-trivial single-qubit Pauli errors that break the code. A distance-$d$ QEC code can detect up to d1d-1 errors and correct up to $\lfloor\frac{d-1}{2}\rfloor$. Therefore, a larger code distance implies a higher capability to correct errors. However, in general, a larger code distance requires a larger number of physical qubits, which is also quantified by the code rate $\frac{k}{n}$. When talking about codes with no provable or unknown distance, as for most VarQEC instances, we simplify the notation to
$((n,K)).$ (16)
The special case of additive QEC codes, usually also called stabilizer codes, is indicated by the modified notation
$[[n,k,d]].$ (17)
@@ 4 Learning Error Correction Codes
In this section, we introduce our method for learning QEC codes by maximizing state distinguishability. We define a loss function based on the trace distance between quantum states, realize the method as a variational quantum algorithm, discuss how to perform error correction using recovery operations, and analyze the properties of the learned codes. Supplementary details can be found in Appendices˜B and C.
@@
[…]
@@ 4.1 Lost Trace and Distinguishability Loss
where $\rho_{L}$ and $\sigma_{L}$ are prepared using the potentially parameterized encoding operation $U_{\mathrm{enc}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\Theta)}$ as described in equation˜10. Due to the trace distance’s invariance under both tensoring with pure states and unitary transformations, we can also express the first term as $T(\rho_{L},\sigma_{L})$, but the former is usually more straightforward to compute. This measure is also indicated in figure˜1 and its components are explicitly marked in figure˜2.
@@
For practical optimization, evaluating the worst-case loss is challenging because it involves a maximization over an infinite set of state pairs. In principle, it is possible to use techniques from semi-definite programming [95], but it is hard to obtain guarantees for the efficiency of finding the solution. Moreover, we empirically found that the worst-case formulation also has the drawback that employing it as a loss function leads to severe instabilities in the training procedure, which aligns with findings from classical machine learning [43, 7]. To address this, we define the average-case distinguishability loss as
$\mathcal{D}(\mathcal{N}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5};\Theta})=\int_{\rho}\int_{\sigma}\Delta_{T}(\rho,\sigma;\mathcal{N}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5};\Theta})\,d\nu(\rho)\,d\nu(\sigma),$ (20)
@@
Conceptually, our distinguishability loss is related to the recoverable quantum information introduced in Ref. [34], as both are built from the trace distance. The recoverable quantum information is defined as $\mathcal{R}_{Q}(t)=\frac{1}{2}\min_{\vec{n}}\left\|\hat{\rho}_{\vec{n}}(t)-\hat{\rho}_{-\vec{n}}(t)\right\|_{1}$, where $\hat{\rho}_{\vec{n}}$ describes the evolution under initial pure state $\vec{n}$, and $\hat{\rho}_{-\vec{n}}$ under the initially orthogonal state, respectively. For the special case of a single logical qubit under isotropic unital noise, which guarantees that the worst-case violation is caused by antipodal pairs, both measures coincide, concretely $\overline{\mathcal{D}}$ is proportional to $1-R_{Q}$. However, the distinguishability loss constitutes a more general and practically different object: It naturally extends to multiple logical qubits per patch, and as shown in section˜4.4 $\overline{\mathcal{D}}$ establishes analytic bounds on the fidelity after recovery, one of the standard figures of merit in QEC. Furthermore, while $\mathcal{R}_{Q}$ is defined as a discrete reward function only, the average-case proxy of the distinguishability loss $\mathcal{D}$ is smooth and differentiable in the encoding parameters, and can therefore be optimized directly with gradient-based techniques. Importantly, the recoverable quantum information assumes unital noise, and the efficient evaluation of $\mathcal{R}_{Q}$ has been tied to either isotropic channels, or more general Pauli noise under restrictions to CHZ circuits [76]. In contrast, the distinguishability loss is defined for arbitrary CPTP maps, and its two-design approximation $\mathcal{D}_{\mathcal{S}}$ ensures efficient evaluation. In this sense, the distinguishability loss generalizes the intuition behind the recoverable quantum information, and establishes an efficiently computable objective for learning noise-tailored encodings under arbitrary noise.
@@ 4.2 Variational Quantum Error Correction
Figure 3: An instance of the REA on 44 qubits. It consists of an initial block of parameterized single-qubit rotations 33 blocks of successive entangling operations. The entangling blocks themselves consist of a CZ gate and a parameterized single-qubit rotation on the two involved qubits. The control and target for each two-qubit block are determined uniformly at random, i.e. the exact instance of an ansatz is defined by the random seed that was used for generating it. A more general version of this ansatz is detailed in figure˜13.
In this work, for a proof-of-principle demonstration of the concept, we choose to formulate the task as a variational quantum algorithm [20]. We first select a variational ansatz $U_{\mathrm{enc}}(\Theta)$ for the encoding unitary, with trainable parameters $\Theta$. Therefore, we utilize a newly developed REA, which features an initial single-qubit layer, followed by randomly placed two-qubit blocks. An exemplary instance of such a layout is shown in figure˜3. The motivation for this ansatz is to allow searching for well-suited shallow circuit instances by iterating over layouts with varying two-qubit placements. Furthermore, a REA can easily be made hardware-efficient by restricting to two-qubit gates that allow for native realization on the hardware. The parameters can be optimized using either gradient-free or gradient-based optimization methods. We employ the quasi-Newton limited-memory BFGS (L-BFGS) algorithm [63] due to its fast convergence for medium-sized models. However, conceptually, other optimization techniques like standard gradient descent can also be used. To enhance performance, in future work, one can also focus on a more structured approach for choosing optimal circuit architectures, using e.g. differentiable architecture search [117] or incrementally growing the ansatz [40].
@@
Figure 4: Pipeline of learning VarQEC encodings with the distinguishability loss (adapted and extended from [70]). Pairs of $k$-qubit states from a two-design $\mathcal{S}$ are prepared and encoded into an $n$-qubit Hilbert space with the encoding unitary $U_{\text{enc}}(\Theta)$. The resulting encoded states are subjected to the noise channel $\mathcal{N}$, before the trace distance between the disturbed states is evaluated. The difference to the baseline trace distance is computed and denoted as the lost trace distance $\Delta_{T}(\rho,\sigma;\mathcal{N};\Theta)$. This information is averaged over all state pairs $\rho,\sigma\in\mathcal{S}$, giving the two-design approximation to the average-case distinguishability loss $\mathcal{D}_{\mathcal{S}}(\mathcal{N};\Theta)$. This objective is minimized with a suitable technique, in this paper instantiated as a variational quantum algorithm utilizing gradient-based optimization. The converged encoding constitutes a $((n,2^{k}))$ code – either for exact or approximate QEC, depending on the setup – finetuned for the noise channel $\mathcal{N}$.
@@
Due to its resemblance to a variational quantum algorithm, we denote this realization of the training procedure as variational quantum error correction. The resulting circuits are dubbed VarQEC codes, and grouped by their code parameters $((n,2^{k}))$. While in section˜4.4 we also derive a methodology to quantify the code distance $d$ of VarQEC codes, in general, and by construction, the VarQEC procedure does not explicitly enforce the ability to recover information under specific error terms. Rather, the distinguishability loss is used to automatically focus on specific error terms to minimize the overall loss of information. The described training procedure is conceptually illustrated in figure˜4.